Given the frequency table below, which of the following is the corresponding relative frequency table?

Value Frequency
2 6
3 4
4 5
5 7
6 3

Select the correct answer below:

a)

Value Frequency
2 0.2
3 0.2
4 0.24
5 0.24
6 0.12

b)

Value Frequency
2 0.24
3 0.16
4 0.24
5 0.24
6 0.12

c)

Value Frequency
2 0.2
3 0.2
4 0.24
5 0.28
6 0.08

d)

Value Frequency
2 0.24
3 0.16
4 0.2
5 0.28
6 0.12
Answer #1

Solution :-

value

Frequency ( F)

Relative F

2

6

6/25 = 0.24

3

4

4/25 = 0.16

4

5

5/25 = 0.2

5

7

7/25 = 0.28

6

3

3/25 = 0.12

Σ F = 25

Correct Option is ' d '

value

Relative ' F'

2

0.24

3

0.16

4

0.2

5

0.28

6

0.12